OpenStudy (anonymous):

Solve, using the substitution method. y = 2x + 6 3x – y = 6

OpenStudy (diyadiya):

y=2x+6 substitute 2x+6 for y 3x- (2x+6) =6 3x-2x-6=6 x-6=6 now add 6 on both sides

OpenStudy (diyadiya):

what do you get for x now?